### v2

# IC-Custom: Diverse Image Customization via In-Context Learning

Abstract
Image customization, a crucial technique for industrial media production, aims to generate content that is consistent with reference images. However, current approaches conventionally separate image customization into position-aware and position-free customization paradigms and lack a universal framework for diverse customization, limiting their applications across various scenarios. To overcome these limitations, we propose IC-Custom , a unified framework that seamlessly integrates position-aware and position-free image customization through in-context learning. IC-Custom concatenates reference images with target images to a polyptych, leveraging DiT’s multi-modal attention mechanism for fine-grained token-level interactions. We introduce the In-context Multi-Modal Attention (ICMA) mechanism with learnable task-oriented register tokens and boundary-aware positional embeddings to enable the model to correctly handle different task types and distinguish various inputs in polyptych configurations. To bridge the data gap, we carefully curated a high-quality dataset of 16K identity-consistent samples with 10K from real-world sources and 6K from high-quality synthetic data, avoiding the overly glossy and over-saturated synthetic appearance. IC-Custom supports various industrial applications, including try-on, accessory placement, furniture arrangement, and creative IP customization. Extensive evaluations on our proposed ProductBench and the publicly available DreamBench demonstrate that IC-Custom significantly outperforms community workflows, closed-source models like GPT-4o, and state-of-the-art open-source approaches. IC-Custom achieves approximately 73% higher human preference across identity consistency, harmonicity, and text alignment metrics, while training only 0.4% of the original model parameters.
Figure 1: Visualization for our IC-Custom results. Our method supports diverse image customization scenarios, including position-aware, position-free, and multi-reference conditions, effectively handling various reference types such as rigid objects, non-rigid objects, anime characters, and real-world photographs.
footnotetext: Corresponding authors
1 Introduction
Image customization, which aims to ensure that generated content maintains a consistent identity with reference images, has brought various applications such as product placement [25, 3, 4], subject-driven generation [36, 42, 48], and visual try-on [44, 10, 46]. These capabilities are particularly significant for industrial media production, enabling consistent content creation across diverse visual contexts.
Early image customization [35, 8, 1] methods relied on per-instance optimization, which was time-consuming. Subsequent methods [47, 4, 3] introduced additional control branches to pre-trained diffusion models to inject identity information from reference images, but these were limited by model architecture constraints and scaling difficulties, leading to suboptimal performance. Recently, leveraging the long-range modeling inductive bias of DiT architectures [29, 7, 19] , image conditions can be directly input as sequences to the model, interacting with noisy tokens via multi-modal attention mechanisms without requiring additional branches, enabling image customization methods to demonstrate powerful emergent capabilities [38, 26, 20, 21, 40].
Despite these advances, existing methods still face significant challenges in maintaining consistent subject identity across diverse user requirements and customization scenarios (see Tab. 1): (1) They typically treat image customization as two distinct tasks: position-aware customization (inserting identity into masked regions of a reference image) and position-free customization (generating identity-consistent images from scratch). (2) They lack comprehensive support for different mask specifications, including both precise masks and user-drawn masks under position-aware customization settings. (3) They do not support multi-reference customization scenarios where the same identity appears across multiple reference images. These limitations restrict the ability to learn comprehensive and robust identity representations, necessitating separate models for each scenario and preventing the development of unified frameworks that can flexibly handle diverse customization requirements.
Table 1: Comparison of IC-Custom with previous image customization methods, FLUX.1-Fill+FLUX.1-Redux workflow [20, 21], OminiCtrl [40], Insert Anything [38], DreamO [26] and GPT-4o [15]. IC-Custom offers flexibility and versatility. The checkmarks and crosses indicate whether each method can handle the corresponding task well.
Model Position-aware Position-free Multi-ref.
precise mask user-drawn mask
FLUX workflow
OminiCtrl
Insert Anything
DreamO
GPT-4o
IC-Custom
To this end, we propose IC-Custom , a unified framework that seamlessly integrates position-aware and position-free image customization, enabling flexible personalization across diverse scenarios. Specifically, we concatenate the reference image as context with the target image (either a partially masked image or a fully masked empty canvas) to form a diptych, which is then input to the model along with mask indicators. Leveraging DiT’s multi-modal attention mechanism, our approach enables the model to transfer identity information from the reference image to the target image during the denoising process (see Fig. 1).
To provide comprehensive customization support across diverse scenarios, we introduce a novel In-context Multi-Modal Attention (ICMA) module that replaces the original Multi-Modal Attention in DiT architectures. This module incorporates two key innovations: (1) Three types of learnable task-oriented register tokens that specify the customization type: position-aware customization (with precise or user-drawn masks) and position-free customization, enabling the model to adaptively adjust its behavior according to distinct user requirements. (2) Three types of learnable positional embeddings that indicate the spatial relationships: condition embeddings (CE) for reference conditions, target embeddings (TE) for the target canvas and index embeddings (IE) to distinguish multiple references. These embeddings help the model clearly differentiate canvas boundaries and eliminate ambiguities in the polyptych format.
To enable effective training of our unified framework, we curated a high-quality dataset consisting of both real-world and synthetic samples. Specifically, we collected 8K identity-consistent diptychs from real-world sources and an additional 4K synthetic diptychs, resulting in a total of 12K diptychs. For multi-reference scenarios, we further collected 2K real-world polyptychs and synthesized 2K polyptychs, yielding a total of 4K polyptychs. This comprehensive dataset enables our model to learn robust identity representations across diverse contexts and viewpoints, while also addressing the limitations of previous methods that overly rely on synthetic data and often produce artificial-looking results.
To extensively evaluate the performance of our method, we utilize ProductBench and DreamBench [35] to assess both position-aware and position-free customization capabilities. ProductBench is our manually collected benchmark for evaluating position-aware customization, comprising 40 identity-consistent images with an even distribution of rigid and non-rigid objects, along with their corresponding precise masks and user-drawn masks. We also use DreamBench to evaluate position-free customization performance. Extensive subjective and objective evaluations demonstrate the effectiveness of our IC-Custom , outperforming community workflows, closed-source GPT-4o, and state-of-the-art open-source approaches. Notably, our method achieves a significant advantage of approximately 73% in human preference ratios across ID consistency, overall harmonicity, and text alignment metrics compared to other methods, while only training 0.3% of the parameters of the original FLUX model.
In summary, our contributions are as follows:
  • We propose a unified framework that seamlessly integrates position-aware and position-free image customization via in-context learning.
  • We introduce a novel ICMA module, which enables the model to handle position-free, position-aware, and multi-reference customization tasks.
  • We curate a dataset from real-world sources, addressing the limitations of existing methods that rely on synthetic data, which often produce artificial-looking results.
  • We demonstrate that our method outperforms existing approaches across a range of metrics, surpassing community workflows, closed-source models, and state-of-the-art open-source methods.
Figure 2: Model overview. 1) Our model receives in-context polyptych inputs and text prompts; 2) It randomly decides whether to mask the entire target canvas to produce position-free or mask-guided position-aware in-context latents; 3) Learnable embeddings that indicate input types are added to the input sequence along with conventional positional encodings before being fed into the DiT. We add LoRA adapters to 20 DiT blocks and unfreeze input layers during training.
2 Related Work
2.1 Image Diffusion Models
Recently, diffusion models [37, 11] have achieved state-of-the-art results in image synthesis quality, outperforming traditional generative models such as Variational Autoencoders (VAE) [16] and Generative Adversarial Networks (GANs) [9] by a large margin. As a result, many current text-to-image methods [6, 11, 27, 31] adopt diffusion models as their image generation framework. Early attempts utilize a U-Net [34] architecture with cross-attention for text-to-image generation, garnering significant attention due to their competitive performance and efficiency. Specifically, the open-sourcing of Stable Diffusion [33] has greatly promoted the development of image synthesis research. More recently, diffusion transformer methods, such as SD3 [7] and Flux [19] model, have achieved superior performance by incorporating transformer architectures [43] into diffusion models. These fundamental text-to-image models have been widely used in many downstream image tasks, including depth estimation, image editing, and more.
2.2 Image Customization
By utilizing basic text-to-image models, image customization can be achieved by incorporating additional conditions during the generation process. One series of work [45, 23, 14, 26, 40, 5] directly generates the target image conditioned on reference images, as seen in GPT-4o [14], DreamO [26], and OminiControl [40]. These methods typically propose a universal framework for position-free image customization. However, these universal models may not perform well under position-aware customization when an additional masked source image is provided as a condition. It is challenging for these universal models to understand and preserve the unedited regions in this masked image. On the other hand, Insert Anything [38] or inpainting-based methods [20] can insert a subject into the full source image or masked source image in a position-aware manner. However, these methods do not support position-free image customization. Concurrent works such as ACE++ [25] and FLUX.1 Kontext [22] share similar ideas with our approach, yet differ in innovative technical details. In this work, we propose a flexible framework that can address both position-aware customization and position-free customization. We also propose a data curation pipeline to collect high-quality real image data from different product images. Benefiting from this framework and high-quality data, our model achieves highly identity consistent customization, which can be used in real production.
3 Method
3.1 Overview
Benefiting from tokenization and long-range multi-modal attention mechanisms, Diffusion Transformers (such as FLUX.1 dev [19]) have demonstrated preliminary zero-shot in-context generation capabilities [13]. Recent studies [40, 41, 26, 23] have leveraged the in-context generation paradigm to construct customization datasets and train corresponding models. Despite impressive advancements, several challenges still remain. For in-context diptych datasets: existing synthetic datasets suffer from inconsistent identity preservation and artificial appearance, while being predominantly constrained to 512×512 resolution that lacks fine-grained details, and the absence of high-quality real-world diptych datasets further exacerbates these limitations. For in-context customization training: existing models lack a unified approach to comprehensively handle mask-guided position-aware customization (with both precise and user-drawn masks), text-guided position-free customization, and multi-reference customization scenarios within a single framework. To tackle these limitations, as shown in Fig. 2, we propose IC-Custom , which features a unified framework for comprehensive customization (Sec. 3.3) and leverages In-Context Multi-Modal Attention (ICMA) to adapt to various customization settings (Sec. 3.4). We also curate a comprehensive customization dataset from both real-world and synthetic sources, with resolutions mostly exceeding 1024×1024 pixels (Sec. 3.5).
3.2 Preliminaries and Background
Contemporary state-of-the-art image customization methods [26, 40] utilize reference images as conditions and fine-tune a conditional rectified-flow model. Typically, these methods first encode the reference image and target image via VAE encoder separately, which are then tokenized into input tokens $\boldsymbol{Z}_{ref}$ and $Z_{0}$. The noisy target tokens $\boldsymbol{Z}_{t}$ are obtained by $\boldsymbol{Z}_{t}=t\boldsymbol{Z}_{0}+\left(1-t\right)\epsilon_{t}$, where $t\in[0,1]$ is the diffusion timestep and $\epsilon_{t}\sim\mathcal{N}(0,\sigma_{t}^{2}I)$ is the Gaussian noise. Subsequently, the reference tokens $\boldsymbol{Z}_{ref}$ and the noisy target tokens $\boldsymbol{Z}_{t}$ are concatenated along the sequence length dimension as the input of DiT during training. The DiT is optimized with the following target image velocity field prediction objective:
$\min_{\theta}\int_{0}^{1}\mathbb{E}\left[\left\|\left(\boldsymbol{Z}_{1}-\boldsymbol{Z}_{0}\right)-\boldsymbol{V}_{\theta}\left([\boldsymbol{Z}_{t},\boldsymbol{Z}_{ref}],\boldsymbol{c_{T}},t\right)\right\|^{2}\right]dt$ (1)
where $[,]$ denotes the concatenation operation along the sequence length dimension, $\boldsymbol{c_{T}}$ refers to the text tokens, $\boldsymbol{V}_{\theta}$ is the velocity field prediction network, the output of $\boldsymbol{V}_{\theta}$ is truncated to the sequence length of $\boldsymbol{Z}_{t}$. Despite remarkable advances, the conditional customization method do not support mask inputs, making it challenging to effectively adapt to position-aware customization applications.
3.3 In-Context Polyptych Customization
In contrast, as shown in the Fig. 2, we propose to employ an unified in-context polyptych  [^1] For brevity, we describe our method in diptych format. framework to support distinct input formats, and design several training strategies that seamlessly integrate diverse customization paradigms. Specifically, we first concatenate the reference image $I_{ref}$ and the target image $I_{0}$ in a diptych format, then encode the concatenated diptych into latents and add noise to form the input $\boldsymbol{X_{t}}$ for the DiT model. The DiT is optimized with a velocity field prediction objective:
$\min_{\theta}\int_{0}^{1}\mathbb{E}\left[\left\|\left(\boldsymbol{X}_{1}-\boldsymbol{X}_{0}\right)-\boldsymbol{V}_{\theta}\left(\tau\left([\boldsymbol{X}_{t};\boldsymbol{X}_{0}^{mask};\boldsymbol{M}]\right),\boldsymbol{c_{T}},t\right))\right\|^{2}\right]dt$
where $[;]$ denotes the concatenation operation along the channel dimension, $\tau$ denotes the tokenization operation, $\boldsymbol{X}_{0}^{mask}$ is the masked image diptych, $\boldsymbol{M}$ is the mask diptych, $\boldsymbol{V}_{\theta}$ represents velocity field predictor.
During training, we strategically apply either global or local masking to the target image with varying probabilities to create the masked image diptych. This approach enables the model to simultaneously develop proficiency in both position-aware and position-free image customization. Notably, position-free customization can be conceptualized as a special case of position-aware customization where the canvas is completely masked (i.e., $\boldsymbol{M}=\boldsymbol{0}$) and the context content is generated entirely according to the input text prompt. Moreover, we selectively provide the model with in-context text prompts, where descriptions of different scenes are composed of [REF-SCENE] and [TARGET-SCENE] components, which helps the model to leverage the context information for the customization in new scenes. Further details are provided in the appendix.
3.4 In-Context Multi-Modal Attention
Although our pipeline seamlessly adapts to various customization settings, it faces several challenges: (1) Task type ambiguity, e.g., the model often incorrectly treats user-drawn masks as precise boundaries, generating content that completely fills and follows the mask shape; (2) Image boundary confusion, as in polyptych settings, the model struggles to differentiate between reference and target regions, resulting in unwanted edge artifacts. To address these issues, we propose ICMA, a variant of multi-modal attention mechanism. As shown in Fig. 3, ICMA integrates two innovative design choices: (1) learnable task-oriented register tokens to specify the customization type (precise masks, user-drawn masks, and position-free), and (2) learnable boundary-aware positional embeddings including Target Embeddings (TE), Condition Embeddings (CE), and Index Embeddings (IE) to specify spatial relationships between reference and target. Formally, the ICMA mechanism operates as follows:
$\text{Pos}(x)$ $=x+[CE+IE;TE]+RoPE(x)$ (3)
$Q$ $=[Pos(Q_{vis});Q_{txt}]$
$K$ $=[Pos(K_{vis});K_{txt};R_{task}^{i}]$
$V$ $=[V_{vis};V_{txt};R_{task}^{i}]$
$h^{\prime}$ $=\text{MHA}(Q,K,V)$ (4)
where $RoPE(\cdot)$ denotes rotary position encoding [39], $Q_{vis}$, $K_{vis}$, $V_{vis}\in\mathbb{R}^{n\times d}$ are the query, key, and value matrices for visual tokens respectively, $Q_{txt}$, $K_{txt}$, $V_{txt}\in\mathbb{R}^{l\times d}$ are the query, key, and value matrices for text tokens respectively, $R_{task}^{i}\in\mathbb{R}^{4\times d}$ represents the $i$-th learnable task-oriented register tokens, and $\text{MHA}(\cdot)$ denotes the Multi-Head Attention operation. Our proposed ICMA module replaces all multi-modal attention in both the double blocks and single blocks of the original MM-DiT architecture.
3.5 In-Context Customization Data Curation
The scarcity of high-quality customization data presents a critical bottleneck in developing robust customization models. Current methods [40, 26, 23] typically rely solely on synthetic data for model training. However, such synthetic data often exhibits limitations in maintaining identity consistency and achieving photorealistic quality, which can compromise the effectiveness of customization models.
Figure 3: In-Context Multi-Modal Attention (ICMA) mechanism. ICMA incorporate several learnable task-oriented register tokens and boundary-aware positional embeddings (including CE, TE, IE) to indicate different task type and input boundaries.
To address this challenge, we introduce ProductData , a high-quality customization dataset designed to ensure both authenticity and diversity. Specifically, we first collect nearly 8K identity-consistent image pairs from e-commerce platforms, covering real-world scenarios such as clothing try-on, cosmetics, furniture, electronics, accessories, home decor, kitchenware, and personal care products, with image resolutions ranging from $800\times 800$ to $3000\times 3664$ pixels. To further enrich the dataset and broaden its coverage beyond commercial products, we supplement these real-world samples with 4K high-quality identity-consistent synthetic pairs, carefully filtered from the SynCD $1024\times 1024$ subset [18]. This results in a comprehensive diptych dataset of 12K samples. For polyptych extension, we additionally curate 2K real-world polyptych samples and synthesize 2K more using our diptych model, thereby providing sufficient diversity for multi-reference customization training.
Our filtering rules include: 1) If the cosine similarity between the DINOv2 [28] features of any two images in an item is less than 0.2, remove the item; 2) If all images in an item are blank backgrounds, remove the item; 3) Ensure that the target image is not a blank background image. These filtering rules enhance identity consistency between paired data and prevent ambiguity in the generation process. Furthermore, we utilize the Qwen-VL2.5 [2] to annotate the captions for our in-context customization training data. We also leverage Grounded SAM [32] to obtain ground truth masks for the same identity, and randomly generate user masks according to rules to support the training of our model.
Table 2: Quantitative comparison on position-aware and position-free image customization. Evaluation results on ProductBench and DreamBench indicate that our IC-Custom consistently support diverse customization tasks and outperforms existing methods across all metrics. Human study results further validate the superiority of our method.
Methods ProductBench DreamBench Human Study
Precise Mask User-drawn Mask DINO $\uparrow$ CLIP $\uparrow$ CLIP-T $\uparrow$ Consistency $\uparrow$ Harmonicity $\uparrow$ Text Alignment $\uparrow$
DINO $\uparrow$ CLIP $\uparrow$ CLIP-T $\uparrow$ DINO $\uparrow$ CLIP $\uparrow$ CLIP-T $\uparrow$
Flux-Fill+Redux [20, 21] 60.80 81.66 31.13 62.26 81.60 31.29 - - - 3.2% 5.3% -
OminiCtrl [26] 57.93 76.06 31.31 - - - 48.29 75.85 36.82 1.5% 2.1% 6.3%
DreamO [26] 62.98 78.86 31.25 - - - 57.69 76.33 36.24 5.4% 3.2% 10.1%
Insert Anything [26] 62.71 81.65 31.24 61.21 81.75 31.44 - - - 6.8% 6.5% -
GPT-4o [15] 61.40 78.53 30.72 62.05 79.87 30.58 54.31 77.38 36.33 4.6% 7.5% 21.4%
IC-Custom (Ours) 63.14 81.92 31.75 63.28 81.95 31.80 65.67 83.19 36.88 78.5% 75.4% 62.2%
4 Experiments
4.1 Experiments Setup
Implementation Details
IC-Custom is built upon a pre-trained text-to-image model FLUX.1-Fill-dev [20]. We train LoRA [12] with rank 64 on the first 10 layers of both single blocks and double blocks, and directly fine-tune the image input layer and text input layer. The total trainable parameters amount to 49.2649.26M, which is only $0.4\%$ of the original FLUX model’s 1212B parameters (consisting of 19 double blocks and 38 single blocks). Compared to previous methods [38, 26] that train LoRA on all layers (e.g., DreamO [26] trained 707M parameters), our approach significantly reduces training costs. The model is optimized on our proposed 12k dataset using the AdamW [24] optimizer with a learning rate of 5e55e-5 and a batch size of 4. We train single-scene reference customization for 20k iterations in the first stage, followed by 10k iterations of multi-reference customization in the second stage.
Data Bucketing
To adapt our model to various input resolutions, we implement data bucketing, ensuring that each batch contains inputs of identical resolution, including ([800, 800], [1024, 1024], [1024, 1280], [1280, 1280], [1504, 1504]). During multi-reference training, we similarly bucket the number of input reference images, ensuring each batch contains the same number of reference images, currently supporting ([1, 2, 3]).
Benchmarks
To assess our model’s performance in both position-aware and position-free customization settings, we utilize our proposed ProductBench and the open-source DreamBench [35] benchmark. ProductBench consists of 40 high-quality identity-consistent items with resolutions exceeding 1024×1024 pixels. Each item contains identity-consistent paired images and corresponding masks, with no overlap with our training data. We utilized SAM [17] to annotate precise masks and manually created user-drawn masks. The dataset is evenly divided into rigid and non-rigid categories, covering various domains including clothing try-on, accessories, bags, furniture, toys, perfume, etc., designed to evaluate position-aware customization. DreamBench comprises 30 items, with each item containing 5-6 identity-consistent images, used to evaluate both position-free customization and multi-reference customization. For single-scene reference, we use the first image as the reference, while for multi-reference, we selected a subset with significant scene variations, where each item contains 3 reference images. Additionally, we utilize Qwen-VL2.5 [2] to generate in-context textual descriptions for the aforementioned benchmarks.
Metrics
Follow established methods [38, 35], we consider 3 objective evaluation metrics across two aspects: identity consistency, and text alignment.
  • Identity Consistency: We calculate the DINO-I Score [28] and CLIP-I [30] Score between reference images and generated images to assess identity preservation.
  • Text Alignment: We employ CLIP-T score [30] to evaluate the model’s ability to follow text prompts.
We also incorporate subjective evaluation metrics: identity consistency, overall harmonicity, and text alignment to assess the performance of customization models.
Figure 4: Qualitative comparison on position-aware image customization. OminiCtrl and DreamO lack mask input support. FLUX.1 workflow produces results with visual artifacts. Insert Anything overfills mask regions in user-drawn scenarios, leading to disharmonious results. In contrast, IC-Custom preserves backgrounds while customizing masked regions, generating high-quality outputs with harmonious lighting, shadows, and perspectives.
Figure 5: Qualitative comparison on position-free image customization. Compared to baselines produce artificial or incorrect results, IC-Custom achieves superior performance in identity preservation, visual harmony, and text alignment. Red circles highlight incorrect regions or details.
4.2 Position-Aware Customization
Quantitative Comparisons
Tab. 2 presents the quantitative comparison between our method and other approaches on ProductBench . We compared our approach with community workflows FLUX.1-Fill with FLUX.1-Redux (FLUX.1 workflow) [20, 21], state-of-the-art DiT-based open-source works OminiCtrl [40], DreamO [26] and Insert Anything [38], as well as the commercial application GPT-4o [14]. It is worth noting that OminiCtrl and DreamO do not support position-aware customization tasks and completely fail to preserve the input background. For these methods, we conducted position-free customization by specifying the background through input text as a suboptimal comparison. We conducted experiments on both precise masks and user-drawn masks. OminiCtrl, trained solely on the synthetic Subject200k [40] dataset, produces images with insufficient quality and identity coherence (low DINO-I, CLIP-I, harmonicity). GPT-4o often alters the background identity (low DINO-I, CLIP-I). For example, in visual try-on tasks, it often regenerates the human subject instead of preserving it. Despite FLUX.1 workflow and Insert Anything being specifically developed for image insertion and serving as a strong baseline for this task, our IC-Custom still achieves the best performance across all metrics on both precise masks and user-drawn masks.
Qualitative Comparisons
In Fig. 4, we show the qualitative comparison between our method and other approaches on ProductBench . OminiCtrl, DreamO and GPT-4o do not support mask input, which significantly limits their ability to perform identity insertion based on masks. FLUX.1 workflow tends to generate results with visual artifacts. Insert Anything tends to completely fill the entire mask area in user-drawn mask scenarios, resulting in visually disharmonious outputs. Thanks to the unified in-context structure and diverse training strategies, IC-Custom can accurately preserve the background and customize the identity in the mask region, generating high-quality images with harmonious lighting, shadows, textures, and materials. More results refer to Fig. 6.
4.3 Position-Free Customization
Quantitative Comparisons
FLUX.1 workflow [20, 21] and Insert Anything [38] do not support position-free customization tasks. Despite our attempts to adapt these methods by masking the entire target canvas to create a position-free customization without background context, they merely copy the input reference and lack the ability to generate novel scenes. Therefore, we exclude these methods from our comparison. GPT-4o [14] demonstrates position-free customization capabilities when provided with modified task instructions. OminiCtrl [40] and DreamO [26], specifically designed for position-free customization, serve as strong competitors in our evaluation. As shown in Tab. 2, we observe that OminiCtrl exhibit poor identity consistency (low DINO-I, CLIP-I), while DreamO and GPT-4o produces images of inferior visual coherence. In contrast, our method, trained in an in-context manner on high-quality real-world data, achieves state-of-the-art performance across all metrics.
Qualitative Comparisons
The qualitative comparison results are shown in Fig. 5. FLUX.1 workflow and Insert Anything methods fail to achieve position-free customization, struggling to generate content that significantly differs from the reference image. OminiCtrl and DreamO tend to produce overly-glossy or over-saturated results, causing the generated content to appear artificial. Despite GPT-4o’s strong instruction-following capabilities, it sometimes loses fine-grained identity details. Comparatively, IC-Custom successfully generates harmonious, identity-consistent, and diverse customized images.
4.4 Multi-Reference Customization
To evaluate multi-reference customization performance, we conducted experiments on a subset of DreamBench by using 1-3 reference images to generate customized results for each target prompt. Multi-reference evaluation is more complex than single-reference, as generated results need to maintain correspondence with multiple reference inputs simultaneously. We employed GPT-4o voting to assess our model along two dimensions: multi-reference identity consistency and text alignment. As shown in Fig. 7, the zero-shot capability of the first-stage diptych model deteriorates significantly as the number of reference images increases, likely due to the challenge of selecting relevant information from multiple references. After multi-reference training, our model effectively preserves richer identity details from multiple references.
4.5 Human Evaluation
We conducted a user study with 20 participants on 50 randomly selected samples from both position-aware and position-free subsets. For each sample, participants were asked to select the best-performing model across three dimensions: ID consistency, harmonicity, and text alignment. As shown in Tab. 2, our method achieves superior human preference over existing approaches
4.6 Ablation Studies
We present ablation studies of our IC-Custom in Tab. 3, examining model architecture, training data sources, and training strategies. We first establish zero-shot performance as our baseline. Our experiments validate several key design choices: 1 Without training the DiT’s image and text input layers (w/o IL), the model struggles to effectively transfer pre-trained diffusion prior for customization tasks, particularly in user-drawn mask scenarios; 2 Training solely on synthetic data (w/o RD) compromises identity consistency and realism; 3 Omitting user-drawn mask data during training (w/o UM) significantly degrades performance when handling imprecise masks. The visual example on the Fig. 8 illustrates these findings—both w/o IL and w/o RD variants produce noticeable artifacts, while w/o UM results in unexpected shape distortions. In contrast, our full model demonstrates superior flexibility and effectiveness.
Table 3: Ablation studies. We evaluate different variants of our model on ProductBench. Zero-shot: zero-shot inference without fine-tuning; w/o IL: without training Input Layers; w/o RD: without using Real Data for training; w/o UM: without using User-drawn Mask for training.
Models Precise Mask User-drawn Mask
DINO-I $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$ DINO-I $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$
Zero-shot 55.49 77.55 31.24 57.63 79.84 31.20
w/o IL 62.00 81.52 31.36 62.13 81.33 31.64
w/o RD 62.38 81.81 31.62 62.71 81.85 31.22
w/o UM 62.65 81.82 31.58 61.30 81.28 31.64
Ours 63.14 81.92 31.75 63.28 81.95 31.80
5 Limitations and Conclusions
This paper proposes IC-Custom , a flexible and effective framework for image customization. Our approach features four key innovations: 1) an In-context customization paradigm, effectively unifying position-free and position-aware image customization; 2) a novel In-context Multi-Modal Attention (ICMA) mechanism to adapt to different customization settings; 3) a dataset of polyptychs sourced from real-world data, along with an evaluation benchmark with balanced distribution of rigid and non-rigid customization. Extensive experiments demonstrate that our IC-Custom achieves state-of-the-art performance in multiple scenarios across identity consistency, overall harmonicity, and text alignment metrics, while fully supporting polyptych configurations.
Despite its achievements, our method still has limitations: it lacks explicit modeling of viewpoint and lighting consistency for the reference identity. We will focus on addressing these limitations in future work.